# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ Abstract.
Anxiety, depression, and suicidality are common mental health sequelae following concussion in youth patients, often exacerbating concussion symptoms and prolonging recovery. Despite the critical need for early detection of these mental health symptoms, clinicians often face challenges in accurately collecting patients’ mental health data and making clinical decision-making in a timely manner. Today’s remote patient monitoring (RPM) technologies offer opportunities to objectively monitor patients’ activities, but they were not specifically designed for youth concussion patients; moreover, the large amount of data collected by RPM technologies may also impose significant workloads on clinicians to keep up with and use the data. To address these gaps, we employed a three-stage study consisting of a formative study, interface design, and design evaluation. We first conducted a formative study through semi-structured interviews with six highly professional concussion clinicians and identified clinicians’ key challenges in remotely collecting patient information and accessing patient treatment compliance. Subsequently, we proposed preliminary clinician-facing interface designs with the integration of AI-based RPM technologies (AI-RPM), followed by design evaluation sessions with highly professional concussion clinicians. Clinicians underscored the value of integrating multi-modal AI-RPM technologies to support clinicians’ decision-making while emphasizing the importance of customizable interfaces with explainability and multiple responsible design considerations.
@@ 1. Introduction
The standard of care (SoC) approaches of concussion clinicians include (1) asking patients to recall their at-home behaviors during clinical visits (Silverberg et al., 2020; Brent and Max, 2017) and (2) instructing patients or parents to complete self-evaluation questionnaires at home, such as Generalized Anxiety Disorder-7 (GAD-7) and Patient Health Questionnaire-9 (PHQ-9) (Carson et al., 2021; Silverberg et al., 2020). However, these approaches are not without limitations: it could be challenging for concussion patients to comprehensively and accurately recall their previous behaviors in a weekly clinic visit (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016); patients may be able to complete questionnaires on a regular basis at home, but concussion clinicians can only access the questionnaire results when patients return to the clinics in person, which still delays clinicians’ decision-making.
@@
Previous work has proposed different Remote Patient Monitoring (RPM) technologies to collect patients’ at-home information in various scenarios (Chung et al., 2016; Olen and Dechert-Crooks, 2017; Barnes et al., 2024; van Ierssel et al., 2023; Toresdahl et al., 2021; Tirosh et al., 2024) For example, mobile health (mHealth) applications have been proposed in which patients can report their physiological measurements to care provider teams in real-time. Wearable technologies, such as the ActiGraph (Yang et al., 2020) and Philips Actiwatch (Donahue et al., 2024), have been explored to measure the daily steps and sleep patterns of adult concussion patients. Despite their relative success in capturing patients’ activity data, wearable-based RPM approaches are not without limitations: mHealth applications could lead to cognitive burdens for patients with low digital literacy and brain injuries (Giebel et al., 2024; Purdy, 2023); extensive amount of time and effort is required to possess the massive amount of multi-modal data collected by wearable devices (Ginsburg et al., 2024; Baig et al., 2017). These limitations could result in low compliance among the patients and delays in accessing collected data from clinicians, which further affect timely diagnosis and intervention.
@@
Recent exploration in the HCI community demonstrates the promising potential of leveraging innovative AI technologies for RPM (AI-RPM)to address the aforementioned limitations of traditional RPM approaches. For instance, conversational agents (CAs) powered by large language models (LLMs) enable asynchronous patient-care provider communication to verbally collect patients’ health conditions through regular health check-in conversations (Yang et al., 2024). LLMs have also been explored to organize and summarize key information from massive RPM data on clinician-facing dashboard (Wu et al., 2024). Meanwhile, another line of research explored using AI prediction algorithms with RPM data to assist clinicians’ decision-making in their clinical workflow (Yin et al., 2024).
@@
Despite the promising opportunities of AI-RPM systems, they could lead to additional difficulties and burdens if not appropriately designed. Recent studies discovered that AI prediction models without faithful and reliable explanations cannot provide helpful information (Antoniadi et al., 2021) or actionable insights to clinicians (Rane et al., 2023; Zhang et al., 2024), lose clinician trust (Zhang et al., 2020; Chanda et al., 2024), and even result in human-AI competition (Zhang et al., 2024) Moreover, in high-stakes, high-risk, and uncertain clinical scenarios, clinicians follow strict clinical guidelines and specifications in their daily workflow. AI systems that are in conflict with clinical specifications cannot be used and even cause potential harm to stakeholders (Romero-Brufau et al., 2020).
@@
In contrast, human-centered AI (HCAI) design guidelines (Shneiderman, 2022; Wang et al., 2023b) underscores the importance of engaging stakeholders in the system design process and understanding stakeholders’ workflow, in particular, the needs and challenges they encounter. By following HCI design guidelines, AI systems can be seamlessly integrated into stakeholders’ workflow where stakeholders can confidently understand, trust, and use AI-powered systems in their current workflow. A number of studies have demonstrated the effectiveness of human-centered AI systems in medical and healthcare scenarios to facilitate the needs of diverse stakeholders, including older adults (Hao et al., 2024), cancer patients (Wu et al., 2024), and clinicians (Wang et al., 2023b; Fogliato et al., 2022; Yang et al., 2019; Zhang et al., 2024). Nevertheless, how to employ HCAI guidelines in the design of AI-RPM systems has not been studies yet, and little is known about the challenges concussion clinicians face in remote monitoring youth patients and how AI-RPM technologies could be of practical benefit.
In this work, our objective is to investigate the current workflow of concussion clinicians with respect to the challenges in remote monitoring and decision making of youth concussion patients. In particular, we focus on how clinicians collect patients’ mental health-related information, what decisions they need to make, and what the difficulties are in their current workflow. In addition, we want to explore the design opportunities of AI-RPM technologies in supporting clinicians’ needs for remote monitoring and decision-making during their practice. In this study, we seek to explore two main research questions (RQs):
  • RQ1: What challenges and needs do concussion clinicians face in remote monitoring of mental health symptoms and decision making with multi-modal RPM data for youth concussion patients?
  • RQ2: What are the design considerations needed for AI-RPM technologies to support concussion clinicians in their workflow?
@@
We approached to address these RQs through a three-stage study design: first, we conducted a formative study of semi-structured interviews with six highly professional concussion clinicians and derived a suite of design considerations of AI-RPM technologies that could support clinicians’ need for remote monitoring and clinical decision-making. We came up with preliminary designs of clinician-facing interfaces with different AI-RPM technologies based on design considerations. Subsequently, we engaged the same group of clinicians in a design interface evaluation study to further collect feedback and insights on the design considerations of different AI-RPM technologies. We consolidate the design feedback and revise the preliminary interface design to come up with a complete clinician-facing system design interface, along with the design considerations, as the final artifacts.
@@
This paper makes the following key contributions: (1) we identified the challenges and needs that concussion clinicians encounter during their clinical practice with youth concussion patients, particularly with respect to patient mental health sequelae; (2) we derive a suite of design considerations and visualization feedback to empower advanced AI-RPM technologies to support remote patient monitoring and clinicians’ decision-making; and (3) we deliver a user-centered interface design of a concussion clinician-oriented dashboard system supported by multiple AI-RPM technologies, where the initial interface design and refinements are grounded in users’ needs and feedback.
@@ 2. Related Work
We review prior research on the negative influence youth have due to the mental health sequelae in post-concussion, AI-RPM technologies in clinical settings, as well as designing AI-powered systems for clinical decision support.
2.1. Challenges Related to Mental Health Sequelae Following Concussion
Mental health sequelae, such as anxiety, depression, and suicidality, are common among youth patients (ages 11–17) following a concussion (Stein et al., 2019; Gornall et al., 2021; Ledoux et al., 2022). Among these youth concussion patients, 70–80% of them experience longer recovery periods, normally more than 28 days (Scorza and Cole, 2019; Cuff et al., 2022). Studies have already shown the importance of integrating the assessment of mental health conditions into clinical follow-up visits (Gornall et al., 2021; Russell et al., 2023). Monitoring youth concussion patients’ mental health-related information enables clinicians to timely identify any abnormalities in the patient’s mental health and make mental health-related decisions, such as referring the patient to a mental health expert for further assessment or intervention. Therefore, early identification of mental health sequelae in the initial stages of concussion recovery is crucial for reducing the risk of worsening mental health issues in youth concussion patients and minimizing the negative impact on their concussion recovery.
@@ 2.1. Challenges Related to Mental Health Sequelae Following Concussion
However, there are challenges when it comes to monitoring youth concussion patients’ mental health. Most youth concussion patients are advised to rest at home after a concussion (Silverberg and Mikolić, 2023), their mental health sequelae-related symptoms, such as anxiety, often manifest in their everyday environments, particularly at home, which makes concussion clinicians difficult to detect within clinical settings. Although clinicians can inquire about mental health sequelae-related symptoms during clinical visits, recall bias (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016) may prevent patients from providing complete and accurate information to concussion clinicians. Moreover, clinical follow-up visits for youth concussion patients often lack consistency. Although it is recommended to conduct follow-ups every three days after a concussion (Agency, 2024), one study showed that actual follow-up rates vary significantly, with the lowest being only 13.2% (Ramsay et al., 2023). Insufficient follow-up prevents concussion clinicians from collecting information to accurately assess the youth patient’s mental health sequelae timely, making it difficult to make mental health-related decisions such as conducting further assessments or making a referral to a mental health expert. Therefore, it is essential to remotely monitor and collect behavioral data from youth concussion patients outside of clinical settings to assist clinicians in the early detection of mental health sequelae.
@@ 2.2. AI-RPM technologies in Clinical Settings
Remote Patient Monitoring (RPM) includes primarily wearable devices and home-use medical equipment that collect patient health data outside of clinical settings (Mendel et al., 2024). These systems transmit monitored patient data to healthcare providers for real-time monitoring and assessment, thereby supporting clinicians in their decision-making (Catalyst, 2018; Baig et al., 2017; of Health et al., 2008). Wearable devices, for example, serve as valuable tools for remotely monitoring patient data. Existing studies have leveraged wearable technology to track and collect key physiological data from concussion patients during their recovery period, such as physical activity levels, for recovery progress assessments (Tirosh et al., 2024). Additionally, researchers have investigated the relationship between post-concussion symptoms and physical activity in young patients. For instance, Yang et al. (2020) utilized ActiGraph, a smartwatch-based device, to monitor patients’ physical activity and collect movement-related data, such as step count.
@@
Furthermore, researchers in the Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) fields have started exploring mobile health (mHealth) technologies, which use mobile devices to collect patient information, facilitate self-management, and enhance remote communication between patients and clinicians (Corwin et al., 2024; Nyapathy and Arriaga, 2019; Salamah et al., 2021; West et al., 2018). For instance, El-Rashidy et al. (2021) indicated that mHealth can improve diagnostic efficiency and reduce healthcare costs for patients. RPM systems can display collected data via user interfaces for patient review (Griggs et al., 2018) or integrate the data into electronic health record (EHR) systems (Dinh-Le et al., 2019). However, given the significant time constraints that clinicians already face (Häikiö et al., 2020), along with their existing workload, the additional data generated by RPM systems could further exacerbate these challenges.
With recent advances in AI, AI-driven tools have the potential to assist clinicians by analyzing and summarizing large volumes of patient data and presenting information in a more interpretable format, thus reducing the cognitive load in clinical workflows (Häikiö et al., 2020). Research suggests that combining subjective patient perceptions with quantifiable metrics enhances the value and interpretability of patient data (Häikiö et al., 2020). For instance, voice assistants powered by large language models (LLMs) can interact with patients using natural language, capture subjective symptom experiences, and visualize them on a dashboard for clinicians (Yang et al., 2024). AI can also be leveraged for predictive analytics to help forecast patient health outcomes (Zhang et al., 2024; Pathak, 2024; Collins and Moons, 2019). Dabek et al. (2022) developed an AI model to predict the risk of mental health sequelae following a concussion, achieving 88.2% accuracy in predicting outcomes within 90 days post-injury. Despite these advancements, clinicians often remain skeptical about the accuracy of AI-generated predictions (Bohr and Memarzadeh, 2020), which often arises from the lack of transparency in AI decision-making. Providing clinicians with a clear understanding of AI’s limitations and actionable insights can enhance their understanding of AI outputs (Liao et al., 2020), thus increasing confidence in AI-assisted decision-making. Recent studies emphasize that AI explanations should align with clinicians’ cognitive models (Sendak et al., 2020; Sivaraman et al., 2023) to enhance trust and acceptance of AI-generated recommendations. Additionally, supporting AI predictions with evidence-based references and citations may further help clinicians validate AI outputs and increase confidence in AI-assisted decision-making. AI has the potential to support remote patient monitoring (RPM) systems by enhancing the efficiency of data collection and processing while supporting clinical decision-making. However, poor system design may bring new challenges and workload for clinicians. Therefore, integrating well-thought-out design with AI-RPM systems is important to ensure that they enhance rather than hinder clinicians’ work efficiency.
2.3. Designing AI-Powered Systems for Clinical Decision Support
In recent years, with the rapid advancement of artificial intelligence (AI) technology, human-computer interaction (HCI) researchers have increasingly explored how AI can better support clinicians in making clinical decisions. Their research spans various medical domains, including cancer diagnosis (Cai et al., 2019b, a; Denekamp et al., 2007), cardiotoxicity detection (Ahmed et al., 2024), diabetic retinopathy screening (Beede et al., 2020), and rehabilitation assessment (Lee et al., 2021, 2020).

### v2

@@ Abstract.
Anxiety, depression, and suicidality are common mental health sequelae following concussion in youth patients, often exacerbating concussion symptoms and prolonging recovery. Despite the critical need for early detection of these mental health symptoms, clinicians often face challenges in accurately collecting patients’ mental health data and making clinical decision-making in a timely manner. Today’s remote patient monitoring (RPM) technologies offer opportunities to objectively monitor patients’ activities, but they were not specifically designed for youth concussion patients; moreover, the large amount of data collected by RPM technologies may also impose significant workloads on clinicians to keep up with and use the data. To address these gaps, we employed a three-stage study consisting of a formative study, interface design, and design evaluation. We first conducted a formative study through semi-structured interviews with six highly professional concussion clinicians and identified clinicians’ key challenges in remotely collecting patient information and accessing patient treatment compliance. Subsequently, we proposed preliminary clinician-facing interface designs with the integration of AI-based RPM technologies (AI-RPM), followed by design evaluation sessions with highly professional concussion clinicians. Clinicians underscored the value of integrating multi-modal AI-RPM technologies to support their decision-making while emphasizing the importance of customizable interfaces through collaborative design and multiple responsible design considerations.
@@
Figure 1. Refined AI-RPM system designed to support concussion clinicians in mental health monitoring: (A) patient list, (B) patient basic information, (C) sleep and physical activities, (D) self-report symptoms, (E) questionnaire result, and (F) AI risk score prediction for mental health sequelae.
@@ 1. Introduction
The standard of care (SoC) approaches of concussion clinicians include: (1) asking patients to recall their at-home behaviors during clinical visits (Silverberg et al., 2020; Brent and Max, 2017), and (2) instructing patients or parents to complete self-evaluation questionnaires at home, such as Generalized Anxiety Disorder-7 (GAD-7) and Patient Health Questionnaire-9 (PHQ-9) (Carson et al., 2021; Silverberg et al., 2020). However, these approaches are not without limitations: it could be challenging for concussion patients to comprehensively and accurately recall their previous behaviors in a clinic visit (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016); patients may be able to complete questionnaires on a regular basis at home, but concussion clinicians can only access the questionnaire results when patients return to the clinics in person, which still delays clinicians’ decision-making.
@@
Previous work has proposed different Remote Patient Monitoring (RPM) technologies to collect patients’ at-home information in various scenarios (Chung et al., 2016; Olen and Dechert-Crooks, 2017; Barnes et al., 2024; van Ierssel et al., 2023; Toresdahl et al., 2021; Tirosh et al., 2024). For example, mobile health (mHealth) applications have been proposed in which patients can report their physiological measurements to care provider teams in real time. Wearable devices, such as the ActiGraph (Yang et al., 2020) and Philips Actiwatch (Donahue et al., 2024), have been explored to measure the daily steps and sleep patterns of adult concussion patients. Despite their relative success in capturing patients’ activity data, these traditional RPM approaches are not without limitations: mHealth applications could lead to cognitive burdens for patients with low digital literacy and brain injuries (Giebel et al., 2024; Purdy, 2023); extensive amount of time and effort is required to process the massive amount of multi-modal data collected by wearable devices (Ginsburg et al., 2024; Baig et al., 2017). These limitations may reduce patient compliance and lead to clinicians delaying collected data access, which further affects timely diagnosis and intervention.
@@
Recent exploration in the HCI community demonstrates the promising potential of leveraging innovative AI technologies for RPM (AI-RPM) to address the aforementioned limitations of traditional RPM approaches. For instance, conversational agents (CAs) powered by large language models (LLMs) enable asynchronous patient-provider communication to verbally collect patients’ health conditions through regular health check-in conversations (Yang et al., 2024). LLMs have also been explored to organize and summarize key information from massive RPM data on clinician-facing dashboards (Wu et al., 2024). Meanwhile, another line of research explored using AI prediction algorithms with RPM data to assist clinicians’ decision-making in their clinical workflow (Yin et al., 2024).
@@
Despite the promising opportunities of AI-RPM systems, they could lead to additional difficulties and burdens if not appropriately designed. Recent studies discovered that AI prediction models without faithful and reliable explanations cannot provide helpful information (Antoniadi et al., 2021) or actionable insights to clinicians (Rane et al., 2023; Zhang et al., 2024), lose clinician trust (Zhang et al., 2020; Chanda et al., 2024), and even result in human-AI competition (Zhang et al., 2024). Moreover, in high-stakes, high-risk, and uncertain clinical scenarios, clinicians follow strict clinical guidelines and specifications in their daily workflow. AI systems misaligned with clinical specifications may be unusable and even cause potential harm to stakeholders (Romero-Brufau et al., 2020).
@@
In contrast, human-centered AI (HCAI) design guidelines (Shneiderman, 2022; Wang et al., 2023b) underscore the importance of engaging stakeholders in the system design process and understanding stakeholders’ workflow, in particular, the needs and challenges they encounter. By following HCAI design guidelines, AI systems can be seamlessly integrated into stakeholders’ workflow where stakeholders can confidently understand, trust, and use AI-powered systems in their current workflow. A number of studies have demonstrated the effectiveness of human-centered AI systems in medical and healthcare scenarios to facilitate the needs of diverse stakeholders, including older adults (Hao et al., 2024), cancer patients (Wu et al., 2024), and clinicians (Wang et al., 2023b; Fogliato et al., 2022; Yang et al., 2019; Zhang et al., 2024). Nevertheless, how to employ HCAI guidelines in the design of AI-RPM systems has not been studied yet, and little is known about the challenges concussion clinicians face in remote monitoring youth patients and how AI-RPM technologies could be of practical benefit.
@@
In this work, our objective is to investigate the current workflow of concussion clinicians with respect to the challenges in remote monitoring and decision-making of young concussion patients. In particular, we focus on how clinicians collect patients’ mental health-related information, what decisions they need to make, and what the difficulties are in their current workflow. In addition, we want to explore the design opportunities of AI-RPM technologies in supporting clinicians’ needs for remote monitoring and decision-making during their practice. To answer these questions, we conducted a formative study of semi-structured interviews with six highly professional concussion clinicians and derived a suite of design considerations of AI-RPM technologies that could support clinicians’ needs and clinical decision-making. We then came up with preliminary system designs based on design considerations. Subsequently, we engaged the same group of clinicians in an evaluation study to further collect feedback and insights on the design. Finally, we consolidate the feedback and revise the design to come up with a complete clinician-facing system design as the final artifact.
@@
This paper makes the following key contributions: (1) we identified the challenges and needs that concussion clinicians encounter during their clinical practice with youth concussion patients, particularly with respect to patient mental health sequelae; (2) we derive a suite of design considerations and visualization feedback to empower advanced AI-RPM technologies to support remote patient monitoring and clinicians’ decision-making; and (3) we deliver a user-centered system design of a concussion clinician-oriented dashboard system supported by multiple AI-RPM technologies, where the preliminary system design and refinements are grounded in users’ needs and feedback.
@@ 2. Related Work
Based on the prior research, we first identified the challenges related to mental health sequelae after a concussion. Then, we provided a brief overview of current AI-RPM technologies in clinical settings, as well as AI-based system design for clinical decision support.
2.1. Challenges in Post-Concussion Mental Health Sequelae
Mental health sequelae, such as anxiety, depression, PTSD, and suicidality, are common among youth concussion patients (ages 11–17) (Stein et al., 2019; Gornall et al., 2021; Ledoux et al., 2022). Yang et al. (Yang et al., 2015) studied how concussions impacted the mental health of 71 collegiate athletes and found that 14 reported symptoms of depression, 24 experienced symptoms of anxiety, and 10 reported both depression and anxiety. Another study showed that youth concussion patients tend to have twice the risk of suicide compared to their peers who have never had a concussion (Fralick et al., 2019). Moreover, mental health sequelae can prolong youth concussion patients’ concussion recovery (Silverberg et al., 2020; Iverson and Lange, 2003), and hinder their ability to attend school and participate in normal social activities (Iadevaia et al., 2015), which can be an important way to maintain mental well-being (Zamfir, 2015). Thus, it is important to integrate the assessment of mental health conditions into clinical follow-up procedures (Gornall et al., 2021; Russell et al., 2023). The assessment enables concussion clinicians to identify abnormalities in youth patient mental health conditions and conduct intervention, such as conducting a further assessment or referring the patient to a mental health expert.
@@ 2.1. Challenges in Post-Concussion Mental Health Sequelae
However, there are challenges when it comes to accessing youth concussion patients’ mental health conditions. Most youth patients are advised to rest at home after a concussion (Silverberg and Mikolić, 2023). Their mental health symptoms, such as anxiety, changes in sleep patterns, and disinhibition (Brent and Max, 2017), often manifest outside of clinics, which makes it difficult for concussion clinicians to observe these symptoms in clinical settings. Although clinicians can inquire about mental health symptoms during clinical visits, recall bias (Frissa et al., 2016; Van den Bergh and Walentynowicz, 2016) may prevent patients from providing complete and accurate information to concussion clinicians. Self-report questionnaires like the GAD-7 and PHQ-9 are available for anxiety and depression screening (Brent and Max, 2017). However, clinicians often do not receive the results until the patient’s next clinical visit. Moreover, clinical follow-up visits for youth concussion patients are often insufficient and lack consistency. Although it is recommended to conduct follow-ups every three days after a concussion (Agency, 2024),  Ramsay et al. (2023) found that actual follow-up rates vary significantly, with the lowest being only 13.2%. Insufficient follow-up prevents concussion clinicians from timely collecting mental health information and assessing the youth patient’s mental health conditions. Accordingly, it’s difficult for concussion clinicians to make timely mental health-related decisions. Therefore, novel approaches are needed to support clinicians in remote monitoring the mental health symptoms of youth concussion patients on time.
@@ 2.2. AI-RPM Technologies in Clinical Settings
RPM includes primarily wearable devices and home-use medical equipment that collect patient health data outside of clinical settings (Mendel et al., 2024). These systems transmit monitored patient data to healthcare providers for real-time monitoring and assessment, thereby supporting clinicians in their decision-making (Catalyst, 2018; Baig et al., 2017; of Health et al., 2008). Wearable devices, for example, serve as valuable tools for remotely monitoring patient data. Existing studies have leveraged wearable technology to track and collect key physiological data from concussion patients during their recovery period, such as physical activity levels, for recovery progress assessments (Tirosh et al., 2024). Additionally, researchers have investigated the relationship between post-concussion symptoms and physical activity in young patients. For instance, Yang et al. (2020) utilized ActiGraph, a smartwatch-based device, to monitor patients’ physical activity and collect movement-related data, such as step count.
@@
Furthermore, researchers in the Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) fields have started exploring mHealth applications that facilitate self-management, and enhance remote communication between patients and clinicians (Corwin et al., 2024; Nyapathy and Arriaga, 2019; Salamah et al., 2021; West et al., 2018). For instance, El-Rashidy et al. (2021) indicated that mHealth can improve diagnostic efficiency and reduce healthcare costs for patients. RPM systems can display collected data via user interfaces for patient review (Griggs et al., 2018) or integrate the data into electronic health record (EHR) systems (Dinh-Le et al., 2019). However, given the significant time constraints that clinicians already face (Häikiö et al., 2020), the additional data generated by RPM systems could further exacerbate these challenges.
Recently, AI-driven tools have shown the potential to assist clinicians in reducing the cognitive load in clinical workflows by analyzing and summarizing large volumes of patient data or by presenting information in a more interpretable format  (Häikiö et al., 2020). For instance, voice assistants powered by LLMs can interact with patients using natural language, capture subjective symptom experiences, and visualize them on a dashboard for clinicians (Yang et al., 2024). Despite AI having the potential to enhance today’s RPM systems, poor system design of such AI-RPM may introduce new challenges; Thus, we aim to explore the best design practices to ensure these systems can enhance rather than hinder clinicians’ work efficiency.
2.3. Designing AI-based Systems for Clinical Decision Making
Recently, HCI researchers have had an increasing interest in the clinical-AI decision making research topic, and their works span various medical domains, such as cancer diagnosis (Cai et al., 2019b, a; Denekamp et al., 2007), cardiotoxicity early detection (Ahmed et al., 2024), diabetic retinopathy screening (Beede et al., 2020), and rehabilitation assessment (Lee et al., 2021, 2020). Among some of these works, AI was leveraged for predictive analytics to help forecast patient health outcomes (Zhang et al., 2024; Pathak, 2024; Collins and Moons, 2019). Dabek et al. (2022) developed an AI model with 88.2% accuracy for predicting the risk of mental health sequelae within 90 days post-concussion.